# Deep Learning-based Target-To-User Association in Integrated Sensing and Communication Systems

### v1

@@ II-B Radar signal model and image synthesis
where (i) the number of targets in the environment is denoted by $Q$; (ii) $\beta^{k}_{q,\ell}$ is the complex scattering amplitude the $q$th target, comprising the path-loss, the reflectivity of the target (radar cross section, RCS) and the target’s proper scattering phase (i.e., the phase modeling the interaction of the EM wave with the target); (iii) $\tau_{q,\ell}$ is the two-way delay from the Tx antenna to the target and back to the Rx antenna, while (iv) $z^{k}_{\ell}(t)\sim\mathcal{CN}(0,\sigma_{z}^{2}\delta_{\ell-m}\delta(t))$ is the additive noise, uncorrelated in time and space (across channels). We denote with $s_{\ell}^{k}(t;\mathbf{p}_{q})$, for convenience, the Rx signal due to the $q$-th target located in $\mathbf{p}_{q}$. The delay is
@@
The model of the scattering amplitude is
$\beta^{k}_{q,\ell}=\sqrt{\frac{(4\pi)^{-4}\lambda_{0}^{2}G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})}{\|\mathbf{p}_{q}-\mathbf{p}_{\mathrm{Tx},\ell}\|^{2}\|\mathbf{p}_{\mathrm{Rx},\ell}-\mathbf{p}_{q}\|^{2}}\Gamma_{q}}\;e^{j(\delta_{q}+2\pi\nu_{q}kT_{P})}$ (9)
where $\lambda_{0}=c/f_{0}$ is the carrier wavelength, $G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})$ is the power gain of the single radar antenna element in the direction of the $q$th target $\boldsymbol{\vartheta}_{q,\ell}$, $\Gamma_{q}$ is the radar cross section (RCS) of the target, $\nu_{q}=2V_{q}/\lambda_{0}$ is the Doppler shift of the $q$th target, the function of its radial velocity $V_{q}$ and $\delta_{q}\sim\mathcal{U}[0,2\pi)$ is an additional random phase term modeling unknown phase terms pertaining to the Tx/Rx circuitry as well as to the target. In (7), the $q$th target manifests with a time-invariant term comprising the propagation phase $2\pi f_{0}\tau_{q,\ell}$ as well as a quadratic phase term w.r.t. the delay $-\pi\mu\tau_{q,\ell}^{2}$ [^3] The quadratic phase term $-\pi\mu\tau_{q,\ell}^{2}$ is sometimes referred to as ”video phase” in the radar jargon, and it usually negligible for short-range radars (up to few tens to a hundred of meters in range) compared to the propagation phase [REF]. and a time-varying linear phase term whose frequency is proportional to the delay $\tau_{q,\ell}$.
@@ IV Proposed method
In this section, we provide an overview of the proposed method for the integration of deep learning-based radar multi-target detection and beam prediction aimed at T2U association. The proposed method is sketched in Fig. 4. To clarify the aim of the depicted pipeline, we here briefly describe the proposed processing steps. The system inputs are radar range-angle (RA) image and a set of analog precoding beamforming vectors $\mathbf{f}$ selected at the BS for each vehicular target in the scene. The two inputs are assumed to be synchronized. In the following, we assume that the targets are vehicles crossing the scenario and that each vehicle is equipped with a VE as described in Section II-A. Hence, we investigate the identification of radar targets within RA images, the distinguishability of the corresponding VEs in the beamspace, and the association of radar targets to VEs. A description of the considered hybrid sub-connected communication architecture is provided in Section II-A, along with a specification of the considered analog beamforming codebook. We refer the reader to Section V for a thorough description of the simulation framework, post-processing steps, and radar RA images generation.

### v2

@@ II-B Radar signal model and image synthesis
where (i) the number of targets in the environment is denoted by $Q$; (ii) $\beta^{k}_{q,\ell}$ is the complex scattering amplitude the $q$th target, comprising the path-loss, the reflectivity of the target (radar cross section, RCS) and the target’s proper scattering phase (i.e., the phase modeling the interaction of the EM wave with the target); (iii) $\tau_{q,\ell}$ is the two-way delay from the Tx antenna to the target and back to the Rx antenna, while (iv) $z^{k}_{\ell}(t)\sim\mathcal{CN}(0,\sigma_{z}^{2}\delta_{\ell-m})$ is the additive white Gaussian noise, uncorrelated in space (across channels). We denote with $s_{\ell}^{k}(t;\mathbf{p}_{q})$, for convenience, the Rx signal due to the $q$-th target located in $\mathbf{p}_{q}$. The delay is
@@
The model of the scattering amplitude is [35]
$\beta^{k}_{q,\ell}=\sqrt{\frac{(4\pi)^{-3}\lambda_{0}^{2}G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})}{\|\mathbf{p}_{q}-\mathbf{p}_{\mathrm{Tx},\ell}\|^{2}\|\mathbf{p}_{\mathrm{Rx},\ell}-\mathbf{p}_{q}\|^{2}}\Gamma_{q}}\;e^{j(\delta_{q}+2\pi\nu_{q}kT_{P})}$ (9)
where (i) $\lambda_{0}=c/f_{0}$ is the carrier wavelength, (ii) $G_{\ell}(\boldsymbol{\vartheta}_{q,\ell})$ is the power gain of the single radar antenna element in the direction of the $q$th target $\boldsymbol{\vartheta}_{q,\ell}$, (iii) $\Gamma_{q}$ is the radar cross section (RCS) of the target, (iv) $\delta_{q}\sim\mathcal{U}[0,2\pi)$ is a time-invariant phase term within the CPI, modeling both the Tx/Rx circuitry as well as the interaction of the incident EM wave with the $q$th target and (v) $2\pi\nu_{q}kT_{P}$ is the linear phase term with the slow time function of the Doppler shift $\nu_{q}=2V_{q}/\lambda_{0}$ (for radial velocity $V_{q}$). As the target’s phase $\delta_{q}$ is not known a-priori, it is customary to model it as an unknown deterministic variable, following a Swerling 0 model [35]. In (7), the $q$th target manifests with a time-invariant term comprising the propagation phase $2\pi f_{0}\tau_{q,\ell}$ as well as a quadratic phase term w.r.t. the delay $-\pi\mu\tau_{q,\ell}^{2}$ [^3] The quadratic phase term $-\pi\mu\tau_{q,\ell}^{2}$ is sometimes referred to as ”video phase” in the radar jargon, and it usually negligible for short-range radars (up to few tens to a hundred of meters in range) compared to the propagation phase. and a time-varying linear phase term whose frequency is proportional to the delay $\tau_{q,\ell}$.
@@ IV Proposed Method
In this section, we provide an overview of the proposed method for the integration of deep learning-based radar multi-target detection and beam prediction aimed at T2U association. Motivated by the state-of-the-art performance of deep learning architectures for pattern recognition on complex structured data, we frame the radar multi-target detection and beam prediction tasks under a data-driven perspective, identifying in the YOLO model a consistent architectural bias to flexibly integrate both tasks. The proposed method is sketched in Fig. 4. To clarify the aim of the depicted pipeline, we here briefly describe the proposed processing steps. The system inputs are radar range-angle (RA) image and a set of analog precoding beamforming vectors $\mathbf{f}$ selected at the BS for each vehicular target in the scene. The two inputs are assumed to be synchronized. In the following, we assume that the targets are vehicles crossing the scenario and that each vehicle is equipped with a VE as described in Section II-A. Hence, we investigate the identification of radar targets within RA images, the distinguishability of the corresponding VEs in the beamspace, and the association of radar targets to VEs. A description of the considered hybrid sub-connected communication architecture is provided in Section II-A, along with a specification of the considered analog beamforming codebook. We refer the reader to Section V for a thorough description of the simulation framework, post-processing steps, and radar RA images generation.
@@ V-C T2U Association
The derivation of a theoretical expression for the probability of correct association and its theoretical relationship in the beamspace with multi-target detection is a complex task when generic 2D input radar images are considered. We believe that, owing to its complexity, it can pertain to a completely new contribution that we consider for prospective research, complementary to the empirical study here proposed. This investigation of the target to user association problem from a data-driven point of view has been motivated by the unparalleled empirical performance of deep learning models on the extraction of meaningful features from complex structured data. Therefore, we focused on the identification of the architectural biases of deep learning models [52] faithfully representing the problem at hand. Owing to the pattern recognition capabilities of the latest deep learning models for multi-target detection and classification [53], and to the flexibility provided by the specification of a loss function to use for the optimization of learning models, we have focused on the empirical evaluation of target-to-user association when radar multi-target detection/classification and beam prediction are considered as end-to-end learned downstream tasks. To pursue a fair empirical evaluation of the proposed method, we designed a realistic simulation framework comprising vehicular simulations and dynamic radar and wireless communication simulations, in order to train and test the proposed models across data considering the presence of dynamic complex vehicle meshes in the environment.